OpenStudy (highschoolmom2010):

A frustum of a cone is the part that remains when the top of the cone is cut off by a plane parallel to the base. A popcorn container 9 in. tall is the frustum of a cone. Its small radius is 4.5 in. and its large radius is 6 in. What is its volume?

OpenStudy (anonymous):

1. find the volume of a cone that doesn't have the top cut off and has a base of 6in radius 2. find the volume of a cone that has a radius of 4.5in base 3. find the difference of volume from step 1 and step 2. this is the volume of the frustum.

OpenStudy (highschoolmom2010):

\[V=\frac{ 1 }{ 3 }\pi (6^2)(9)\] \[V=\frac{ 1 }{3 }\pi (36)(9)\]\[V=\frac{ 1 }{3 }\pi (324)\] \[V= (108)\pi\] \[V=339.29\] for cone with no top cut off

OpenStudy (anonymous):

is the cone 9 inches tall or just the popcorn container?

OpenStudy (highschoolmom2010):

it says the container but i dont get how to do it any other way

OpenStudy (anonymous):



OpenStudy (highschoolmom2010):

oh ok

OpenStudy (anonymous):

 h = 9+y use similar triangles: h:6 as y:4.5 or h/6 = y/4.5 replace h with the eq 9+y (9+y)/6 = y /4.5 solve for y

OpenStudy (highschoolmom2010):

ok

OpenStudy (anonymous):

*hint y will be near 25

OpenStudy (anonymous):

i gotta go, but i'll give you some hints/help: remember how I wrote this: 1. find the volume of a cone that doesn't have the top cut off and has a base of 6in radius 2. find the volume of a cone that has a radius of 4.5in base 3. find the difference of volume from step 1 and step 2. this is the volume of the frustum. for step 1, use r=6 and h = 9+27 for step 2, use r=4.5 and h=27